OpenStudy (anonymous):

How do i do this o.O Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. a.) y - 1 = 4(x - 3) b.) y + 1 = 4(x + 3) c.) y - 3 = 4(x + 1) d.) y + 3 = 4(x - 1)

OpenStudy (anonymous):

http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (anonymous):

I can break this down into steps for you to follow along with if you're confused @ForevernAlways

OpenStudy (anonymous):

@agentx5 will you please :/ like i said, im horrible in this area in math.

OpenStudy (anonymous):

Step 1: Average slope: \[m=\frac{rise}{run}=\frac{\triangle y}{\triangle x}=\frac{y_2-y_1}{x_2-x_1}\] rise being the change in the vertical, run being the change in the horizontal. Think of stair steps.

OpenStudy (anonymous):

okay, o.O so how do i start it?

OpenStudy (anonymous):

Step 2: Re-arrange this into your so-called "point-slope formula" Average slope: \[m=\frac{\triangle y}{\triangle x}=\frac{y_2-y_1}{x_2-x_1}\] \[m=\frac{\triangle y}{\triangle x}\] \[m * \triangle x=\triangle y\] \[m(x_2-x_1)=(y_2-y_1)\] Remember that \(y_2-y_1 = \triangle y\) and \(x_2-x_1 = \triangle x\), we're just substituting here

OpenStudy (anonymous):

Step #3 So now that we got it in a nice form, we'll simply input the slope for m: \[(4)(x_2-x_1)=(y_2-y_1)\]

OpenStudy (anonymous):

okay whats up with the triangles? And how can u use, y2 - y1 equation when you dont have 2 parenthesis?

OpenStudy (anonymous):

The triangle is just a symbol that you should get used to seeing, it has a special meaning: the difference between two things, "final minus initial" :-) You'll see this over, and over, and over in math, chemistry, biology, and physics.

OpenStudy (anonymous):

ahhh i hate this :/ lol i think im guna give up because i dont understand anything your saying :(

OpenStudy (anonymous):

\(y_{final}-y_{initial}\) is "the difference between the final and inital for the y variable". "Delta y" or \(\triangle y\) I'm treating it as if it's own variable. You can use "rise", "delta Y", "y final minus y inital", it all means the SAME thing. Like saying red and rouge, English or french it's the same thing. Don't stress the little things, you're doing fine :D

OpenStudy (anonymous):

well im sorry, im stressing haha. it's too complicated. Im going to 10th grade. And im doing geomotry this yr Dx and i cant get thru these stupid algebra probelm tht ur explaining at all :'(

OpenStudy (anonymous):

Step #4: Pick apart your coordinate you were given @ForevernAlways ( \(x_1\) , \(y_1\) ) ( -3 , -1 ) so... \(x_1\) = ? \(y_1\) = ?

OpenStudy (anonymous):

Do you have something you can write my steps down on? Take notes on?

OpenStudy (anonymous):

I'm a visual, spatial, and hands-on type of learner myself. Symbols are (believe it or not) a weakness of mine.

OpenStudy (anonymous):

yes these symbols are confusing the heck outta me. And its way easier for me to learn person to person. doing this virtual skwl thing im doin is flunkn me behind. But yeah im taking down these steps.

OpenStudy (anonymous):

This is the kind of thing I ask questions on: http://openstudy.com/users/agentx5#/updates/4ff38b41e4b01c7be8c738fa It's been many years of practice to get to this level, more than I want to think about tbh lol

OpenStudy (anonymous):

So believe me, if I can do it, you can to. :-D It just takes time and perseverance.

OpenStudy (anonymous):

And there are lots of helpful, patient people here willing to help for nothing more than a virtual medal here or there :-D

OpenStudy (anonymous):

lol okay. And i may be sounding really confusing, but is there any way you can explain this in simplier terms? Like using the question on top?

OpenStudy (anonymous):

Let's see where you're stuck. You have the formula I gave you in step #1, then in step #2 you use Algebra to change it's shape into a form that is useful. :-) Yes? Multiply both sides by "run", "\(\triangle x\)", or "x\(_2\)-x\(_1\)" Do you see how this works?

OpenStudy (anonymous):

kinda o.O

OpenStudy (anonymous):